OpenStudy (anonymous):

The value of y varies directly with x. What is k the constant of variation for this function, and what is y for the given value of x? y=-1/2 when x =6. find y when x=-3.

OpenStudy (anonymous):

A. K= 1/12 ; -4 B. K= 12 ; -4 C. K= -12 ; 4 D. K= -1/12 ; 1/4

OpenStudy (anonymous):

Use y=kx to find k -1/2=6k

OpenStudy (anonymous):

would that be the answer?

OpenStudy (anonymous):

No, you have to solve for k....